# Enhancement of Weak Interactions in Phase Transitions

### v1

@@ Abstract
Weak interactions cause small parity-violating energy differences between left- and right-handed chiral systems. Although normally tiny, these effects may be significantly enhanced during collective phenomena such as phase transitions. We propose a theoretical model describing the enhancement of weak interactions in phase transitions. The enhancement factor is proportional to the critical number of atoms, $N_{c}$, in the nucleus of the new phase. After the nucleus reaches its critical size, it grows until it fills the entire system. Measurement of the ratio of produced left and right chiral structures may provide a way to measure this critical number $N_{c}$. Experiments where definite spin-chiral structures are formed during a phase transition in crossed electric and magnetic fields, indicate $N_{c}\sim 10^{9}-10^{10}$. An open question is whether a similar enhancement could operate during cosmological phase transitions - thereby boosting CP-violating effects sufficiently to contribute to the observed matter-antimatter asymmetry (baryogenesis).
@@ I Introduction
A number of publications have concluded (see, e.g., Refs. [1, 2, 3, 4, 5]) that the Standard Model gives a baryon-to-photon ratio several orders of magnitude smaller than observations. In this situation it is important to check that we have not overlooked any enhancement mechanism which can play a role in electroweak baryogenesis. Indeed, systems are very sensitive to weak fields during phase transitions. A very interesting experiment was performed in Ref. [6]. One may interpret its result as a nine-orders-of-magnitude enhancement of the effects of small perturbations in phase transitions in condensed matter (see below). If a similar enhancement of CP violation existed in the electroweak phase transition (which presumably created the baryon asymmetry of the Universe), CP violation in the Standard Model might produce a baryon-to-photon ratio compatible with observations. Even if the enhancement is not large enough in the Standard Model, it may be important for other models of baryogenesis involving a phase transition.
@@ II Nucleation Model and Enhancement Factor
In the nucleation picture, fluctuations create small droplets of a new phase. Droplets smaller than a critical size $r_{cr}$ quickly disappear, while larger ones grow to fill the system. The probability of forming a critical nucleus is [30]
@@
where $W_{min}$ is the work required to form a critical nucleus. Using formulas presented in the book [30], it is easy to express $W_{min}$ in terms of the chemical potentials per particle of the old and new phases, $\mu_{1}$ and $\mu_{2}$:
@@
$P=-\frac{W_{min}^{+}-W_{min}^{-}}{2kT}\approx-\frac{N_{c}\Delta E}{2kT},$ (6)
@@ IV Possible implications for baryogenesis
In a static case energy difference between matter and antimatter requires CPT violation. However, CP violation may be sufficient in the non-stationary background of evolving Universe.
For example, in electroweak baryogenesis, bubble walls separate the high-temperature symmetric phase from the low-temperature broken-symmetry phase. CP-violating interactions modify the transmission coefficients $T_{\pm}$ for particles and antiparticles crossing these walls [2]. The resulting differences in reflection and transmission affect both the internal pressures and the critical bubble sizes, thereby producing an asymmetry in bubble nucleation rates. This mechanism acts as a kind of “natural selection,” favoring bubbles containing matter over those containing antimatter.

### v2

@@ Abstract
Weak interactions cause small parity-violating energy differences between left- and right-handed chiral systems. Although normally tiny, these effects may be significantly enhanced during collective phenomena such as phase transitions. We propose a theoretical model describing the enhancement of weak interactions in phase transitions. The enhancement factor is proportional to the critical number of atoms, $N_{c}$, in the nucleus of the new phase. After the nucleus reaches its critical size, it grows until it fills the entire system. Measurement of the ratio of produced left and right chiral structures may provide a way to measure this critical number $N_{c}$. Experiments where definite spin-chiral structures are formed during a phase transition in crossed electric and magnetic fields, indicate $N_{c}\sim 10^{9}-10^{10}$. An open question is whether a similar enhancement could operate during cosmological phase transitions - thereby boosting CP-violating effects sufficiently to contribute to the observed baryon-to-photon ratio.
@@ I Introduction
As known, the Standard Model gives a baryon-to-photon ratio several orders of magnitude smaller than observations. In this situation it is important to check that we have not overlooked any enhancement mechanism which can play a role in electroweak baryogenesis. Indeed, systems are very sensitive to weak fields during phase transitions. A very interesting experiment was performed in Ref. [1]. One may interpret its result as a nine-orders-of-magnitude enhancement of the effects of small perturbations in phase transitions in condensed matter (see below). If a similar enhancement of CP violation existed in the electroweak phase transition (which presumably created the baryon asymmetry of the Universe), CP violation in the Standard Model might produce a baryon-to-photon ratio compatible with observations. Even if the enhancement is not large enough in the Standard Model, it may be important for other models of baryogenesis involving a phase transition.
@@ II Nucleation Model and Enhancement Factor
In the nucleation picture, fluctuations create small droplets of a new phase. Droplets smaller than a critical size $r_{cr}$ quickly disappear, while larger ones grow to fill the system. The probability of forming a nucleus of radius $r$ is [25]
@@
where $W_{min}$ is the minimal work required to form a nucleus, which may be presented as a sum of the negative volume term, reflecting advantage of the new phase, and positive surface term: $W_{min}(r)=-B\,4\pi r^{3}/3+\alpha_{t}\,4\pi r^{2}$. This function has maximum at the critical size $r_{c}=2\alpha_{t}/B$. This looks like $W_{min}(r)$ produces a potential barrier for formation of the nucleus bigger than the critical size, with $r>r_{c}$.
Using formulas presented in the book [25], it is easy to express $W_{min}(r_{c})$ in terms of the chemical potentials per particle of the old and new phases, $\mu_{1}$ and $\mu_{2}$:
@@
$P\sim-\frac{N_{c}\Delta E}{2kT}.$ (6)
@@
Here we see the potential for large enhancement: normally in statistical physics energies per degree of freedom are compared to $kT$, but a critical nucleus may contain billions of degrees of freedom, and the total difference $W_{\min}^{+}-W_{\min}^{-}$ is compared with $kT$.
@@ IV Possible implications for baryogenesis
In a static case, the energy difference between matter and antimatter requires CPT violation. However, CP violation may be sufficient in the non-stationary background of evolving Universe.
For example, in electroweak baryogenesis, bubble walls separate the high-temperature symmetric phase from the low-temperature broken-symmetry phase. CP-violating interactions modify the transmission coefficients $T_{\pm}$ for particles and antiparticles crossing these walls (see, e.g., Refs. [28, 29, 30, 31]). According to Ref. [29], the relative difference in the transmission coefficients for quarks and antiquarks is $\Delta\sim 10^{-4}$ in the Standard Model, while the results of Refs. [30, 31] are significantly smaller. This effect leads to different numbers of particles and antiparticles inside the bubble; in the symmetric phase outside the bubble, rapid equilibration occurs due to sphaleron processes.
The resulting differences in reflection and transmission affect both the internal pressures and the critical bubble sizes, thereby producing an asymmetry in bubble nucleation rates. Indeed, the pressure is generated by reflection, therefore it is reduced by transmission: $P_{\pm}=(1-T_{\pm})P_{0}$, where $P_{0}$ is the pressure for $T_{\pm}=0$. The pressure difference implies a difference in the minimal work,
$d(W_{\min}^{+}-W_{\min}^{-})\sim(P_{+}-P_{-})\,dV.$ (10)
Another effect is the difference in critical sizes $r_{\pm}$ of bubbles containing particles and antiparticles. The mechanical equilibrium condition depends on the pressure inside and outside the bubble and on the radius $r$:
$P_{\mathrm{in}}+B=P_{\mathrm{out}}+\frac{2\alpha_{t}}{r},$ (11)
where $B$ is the (volume) bag constant (from the Higgs-field energy) and $\alpha_{t}$ is the surface tension. Therefore a pressure difference between particle and antiparticle cases produces a difference in the critical radius, giving an additional contribution to $W_{\min}^{+}-W_{\min}^{-}$. These two effects may lead to a “natural selection”: critical bubbles containing matter appear more frequently than those containing antimatter.
@@
An open question is whether the nucleation framework and its $N_{c}$-based enhancement can be extended to a crossover rather than a sharp first-order transition. Even in a crossover, droplet-like regions (”bubbles”) of the emergent phase appear within the symmetric phase; see, e.g., [32] and references therein.